# MirrorCalib: Utilizing Human Pose Information for Mirror-based Virtual Camera Calibration

### v2

@@ V-F A Case Study on 3D Human Pose Estimation
VI Conclusion
This paper introduced the novel task of estimating the virtual camera pose relative to the original camera in exercise videos. To achieve this, a modified eight-point algorithm that incorporates mirror characteristics was proposed and the camera parameters were optimized using prior knowledge of the human body. We enhanced the accuracy of our approach with a RANSAC algorithm to reject all outliers. The effectiveness of our method was demonstrated on synthetic and real datasets and a case study of 3D human pose estimation was also provided to confirm its practical applicability. The optimization process and RANSAC algorithm have the potential to be extended to general multi-view camera pose estimation with humans present in the scene. In future work, we will focus on improving the computational efficiency of MirrorCalib.
[Reflective Essential Matrix Decomposition]
Figure 7: Epipolar geometry of the real and virtual cameras.
When the real and virtual cameras capture different views of a 3D scene in Figure 7, any point in the scene and the centers of the camera must lie on the same plane. This gives us the constraint:
$\overrightarrow{O^{\prime}P}\cdot(\overrightarrow{O^{\prime}O}\times\overrightarrow{OP})=0$ (30)
We can denote the coordinates of point $p$ under the real camera frame with origin $O$ to be $[x,y,z]^{T}$ and the coordinates under the right camera frame with origin $O^{\prime}$ to be $[x^{\prime},y^{\prime},z^{\prime}]^{T}$. Subsequently, we get the equation:
$\begin{bmatrix}x^{\prime}\\ y^{\prime}\\ z^{\prime}\end{bmatrix}^{\mathsf{T}}\left(\mathbf{D}\mathbf{t}\times(\mathbf{D}\mathbf{R}\begin{bmatrix}x\\ y\\ z\end{bmatrix})\right)=0$ (31)
where
$\mathbf{D}=\begin{bmatrix}-1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$ (32)
Accounting for the projection geometry and using $\mathbf{K}$ to denote the intrinsic matrix, the following restrictions can be derived:
$\begin{bmatrix}x\\ y\\ z\end{bmatrix}=\mathbf{z}\mathbf{K}^{-1}\begin{bmatrix}u\\ v\\ 1\end{bmatrix},\begin{bmatrix}x^{\prime}\\ y^{\prime}\\ z^{\prime}\end{bmatrix}=\mathbf{z^{\prime}}\mathbf{K}^{-1}\begin{bmatrix}u^{\prime}\\ v^{\prime}\\ 1\end{bmatrix},$ (33)
$\begin{bmatrix}u^{\prime}\\ v^{\prime}\\ 1\end{bmatrix}\mathbf{K}^{-T}\mathbf{E}\mathbf{K}^{-1}\begin{bmatrix}u\\ v\\ 1\end{bmatrix}=0$ (34)
After we have got the fundamental matrix $\mathbf{F}$, we can obtain the essential matrix by $\mathbf{E}=\mathbf{K}^{\mathsf{T}}\mathbf{F}\mathbf{K}$. Let $\mathbf{t^{\prime}}$ be written as:
$\mathbf{t^{\prime}}=\mathbf{D}\mathbf{t}$ (35)
Then the Essential matrix is as follows:
$\mathbf{E}=[\mathbf{t^{\prime}}]_{\times}\mathbf{D}\mathbf{R}$ (36)
where $[\mathbf{t^{\prime}}]_{\times}$ is a skew-symmetric matrix and can be written as:
$[\mathbf{t^{\prime}}]_{\times}=\mathbf{U}diag(1,1,0)\mathbf{W}^{\mathsf{T}}\mathbf{U}^{\mathsf{T}},\mathbf{W}=\begin{bmatrix}0&-1&0\\ 1&0&0\\ 0&0&1\end{bmatrix}$ (37)
or
$[\mathbf{t^{\prime}}]_{\times}=-\mathbf{U}diag(1,1,0)\mathbf{W}\mathbf{U}^{\mathsf{T}},$ (38)
$\mathbf{U}$ is an orthogonal matrix in Equations (31) and (32). And according to Equations (30), (31), and (32), there are two possibilities for the essential matrix:
$\mathbf{E}$ $=\mathbf{U}diag(1,1,0)\mathbf{W}^{\mathsf{T}}\mathbf{U}^{\mathsf{T}}\begin{bmatrix}-1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}\mathbf{R}$ (39)
$=\mathbf{U}diag(1,1,0)\mathbf{V}^{\mathsf{T}}_{1}$
and:
$\mathbf{E}$ $=\mathbf{U}diag(1,1,0)\left(-\mathbf{W}\mathbf{U}^{\mathsf{T}}\begin{bmatrix}-1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}\mathbf{R}\right)$ (40)
$=\mathbf{U}diag(1,1,0)\mathbf{V}^{\mathsf{T}}_{2}$
Note that the matrices $\mathbf{V}^{\mathsf{T}}_{1}$ and $\mathbf{V}^{\mathsf{T}}_{2}$ are orthogonal. If we observe Equations (33) and (34), it reminds us of singular value decomposition. After singular value decomposition, from Equation (33), we can obtain the rotation matrix by:
$\mathbf{R}_{1}=\begin{bmatrix}-1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}\mathbf{U}\mathbf{W}\mathbf{V}^{\mathsf{T}}$ (41)
or:
$\mathbf{R}_{2}=-\begin{bmatrix}-1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}\mathbf{U}\mathbf{W}^{\mathsf{T}}\mathbf{V}^{\mathsf{T}}$ (42)
And translation $[\mathbf{t^{\prime}}]_{\times}$ can be obtained by:
$[\mathbf{t^{\prime}}]_{\times}=\mathbf{U}diag(1,1,0)\mathbf{W}^{\mathsf{T}}\mathbf{U}^{\mathsf{T}}=-\mathbf{U}diag(1,1,0)\mathbf{W}\mathbf{U}^{\mathsf{T}}$ (43)
Because $\mathbf{E}$ and $-\mathbf{E}$ both satisfy Equation (28), $-[\mathbf{t^{\prime}}]_{\times}$ is also a solution. In addition, we can easily obtain $\mathbf{t}$ from $[\mathbf{t^{\prime}}]_{\times}$. Note that there exist four possible combinations of $\mathbf{R}$ and $\mathbf{t}$ that can yield a solution. However, the correct combination can be determined by verifying whether the reconstructed points are positioned behind the cameras after triangulation. This decomposition algorithm is modified according to our scenario from a previous study [22].

### v3

@@ 4.5 A Case Study on 3D Human Pose Estimation
5 Conclusion
This paper introduced the novel task of estimating the virtual camera pose relative to the original camera in exercise videos. To achieve this, a modified eight-point algorithm that incorporates mirror characteristics was proposed and the camera parameters were optimized using prior knowledge of the human body. We enhanced the accuracy of our approach with a RANSAC algorithm to reject all outliers. The effectiveness of our method was demonstrated on a collected real-world dataset and a case study of 3D human pose estimation was also provided to confirm its practical applicability. The optimization process and RANSAC algorithm have the potential to be extended to general multi-view camera pose estimation with humans present in the scene. In future work, we will focus on improving the computational efficiency of MirrorCalib.